OpenStudy (anonymous):

sin 4 theta hey, how do you simplify sin 4 theta? using double angle theory?

OpenStudy (mathteacher1729):

\[\sin(4\theta) = \sin(2\cdot2\theta)\] You would use the double angle formula twice here.

OpenStudy (anonymous):

sin(4theta)=2sin(2theta)cos(2theta)=4sin(theta)cos(theta)cos(2theta)

OpenStudy (anonymous):

are u sure?

OpenStudy (anonymous):

yes, it is by formula sin(2x)=2sin(x)cos(x)

OpenStudy (anonymous):

ok, thank you. i'll give you a medal then. :)

OpenStudy (anonymous):

thx~

OpenStudy (anonymous):

thank you for helping me.